Draw all possible sample of size 2 from the set A={2,3,4,6,8} with replacement and find their mean values. Now construct a frequency distribution of these mean values. then find the mean of this frequency distribution. then finally verify that this mean is equal to the mean of the set A.


1
Expert's answer
2021-12-28T15:44:39-0500

When drawn with replacement, the total number of combinations that can be obtained is "n^r," where "r" is the number of elements chosen and "n" the number of elements to Choose from.

We have "5^2=25" combinations.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample& Sample\ mean, \bar{x} \\ \hline
 2,2 & 2 \\
 2,3 & 2.5 \\
 2,4 & 3 \\
 2,6 & 4 \\
 2,8 & 5 \\
 3,2 & 2.5 \\
 3,3 & 3 \\
 3,4 & 3.5 \\
 3,6 & 4.5 \\
 3,8 & 5.5 \\
 4,2 & 3 \\
 4,3 & 3.5 \\
 4,4 & 4 \\
 4,6 & 5 \\
4,8 & 6 \\
6,2 & 4 \\
 6,3 & 4.5 \\
 6,4 & 5 \\
 6,6 & 6 \\
6,8 & 7 \\
8,2 & 5 \\
 8,3 & 5.5 \\
 8,4 & 6 \\
 8,6 & 7 \\
8,8 & 8 \\
\end{array}"

Make a probability distribution of the sample means


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 \bar{x} & f & probability \\ 
 2 & 1 & 1/25=0.04 \\
 2.5 & 2 & 2/25=0.08 \\
 3 & 3 & 3/25=0.12 \\
 3.5 & 2 & 2/25=0.08 \\
 4 & 3 & 3/25=0.12 \\
 4.5 & 2 & 2/25=0.08 \\
 5 & 4 & 4/25=0.16 \\
 5.5 & 2 & 2/25=0.08 \\
 6 & 3 & 3/25=0.12 \\
 7 & 2 & 2/25=0.08 \\
 8 & 1 & 1/25=0.04 \\
\end{array}"

"\mu_{\bar{x}}=2(0.04)+2.5(0.08)+3(0.12)+3.5(0.08)"

"+4(0.12)+4.5(0.08)+5(0.16)+5.5(0.08)"

"+6(0.12)+7(0.08)+8(0.04)=4.6"


"\mu=\dfrac{2+3+4+6+8}{5}=4.6"

The mean of the frequency distribution of mean values is equal to the mean of the set A


"\mu_{\bar{x}}=4.6=\mu"